OpenStudy (anonymous):

4 - 2 + 1 + (-1/2)+ (1/4) + (-1/8)... series converging or diverging?? i dont know how to use the tests if theres no actual formula..help please

OpenStudy (anonymous):

group the first and second term, likewise 3rd and 4th...and so on. u'll end up gettin a converging gp sequence....so the sequence is converging.

OpenStudy (anonymous):

ok, but i need to find the sum too so how do i take/find the limit?

OpenStudy (anonymous):

2+1/2+1/8+1/32+...=2(1+1/4+1/16+1/64...)=2/(1-1/4)=8/3

OpenStudy (anonymous):

its a geomentric progression...so use the formula for summing up a converging GP

OpenStudy (anonymous):

hmm ok, ill try to figure that out! thank you